Status Report

NASA ISS On-Orbit Status 6 March 2009

By SpaceRef Editor
March 8, 2009
Filed under , , ,
NASA ISS On-Orbit Status 6 March 2009

All ISS systems continue to function nominally, except those noted previously or below.

CDR Fincke started the day with the daily download of the accumulated data of the experiment SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight) from his Actiwatch to the HRF-1 (Human Research Facility 1) laptop as part of another week-long session with SLEEP, his third. [To monitor the crewmember’s sleep/wake patterns and light exposure, the crewmembers wear a special Actiwatch device which measures the light levels encountered by them as well as their patterns of sleep and activity throughout the Expedition and use the payload software for data logging and filling in questionnaire entries in the experiment’s laptop session file on the HRF-1 laptop. The log entries are done within 15 minutes of final awakening for seven consecutive days, as part of the crew’s discretionary “job jar” task list.]

Also before breakfast, Fincke started the next part (3rd of 5) of the periodic acoustic measurement protocol by recording post-sleep data of the crew-worn acoustic dosimeters, later deploying the dosimeters statically (Part 4), one at the SM (Service Module) Central Post, one in Node-2 and the third in an empty rack bay in the Kibo JPM (JEM Pressurized Module), away from air flow. Afterwards (~4:50pm EST), Fincke will record the data taken by the three static dosimeters during the day (Part 5). [Acoustic data must be taken twice per Increment, each time for the duration of the 16-hour crew workday.]

For the Russian Orlan EVA-21A, the CDR gathered the US equipment to be used on the spacewalk.

Mike Fincke & Yuri Lonchakov continued EVA preparations in the DC1 (Docking Compartment), including –

  • Leak checking & valve testing on Orlan and BSS (Orlan Interface Unit) for both suits (#26, #27);
  • Setting up Orlan BRTK “Korona” comm configuration for both suits, then running voice checks & testing medical parameter acquisition of the BETA-08 ECG (electrocardiograph) harnesses with the “Gamma-1M” med complex from the PKO med exam panel for vital signs & equipment monitoring;
  • Mounting the Fresnel lens viewing aid in the helmets;
  • Filling the DIDBs (disposable in-suit drink bags) and installing them in the suits;
  • Configuring the auxiliary NASA equipment collected by Mike for use in the Orlans (no helmet lights this time), plus
  • Taking photos of the outfitted Orlans for downlink.

In addition, Lonchakov wrapped up the re-work of the European EXPOSE-R payload, getting it ready for its external installation. [The failed bolt between EXPOSE-R frame and the external structure that would not engage during EVA-21 was replaced, as were the other three bolts between the sample trays and EXPOSE-R. The container lid and exposed bolts were secured with Aramide tape for safety during egress and translation to its external installation site. NOTE: The spacewalkers were cautioned to take great care in preventing any inadvertent jettisoning/loss of objects during the EVA which could pose conjunction danger to the Shuttle after its launch on the next day and rendezvous on 3/13. More analysis is underway.]

In the US Airlock, FE-2 Magnus meanwhile continued preparations for the 15A spacewalks by –

  • Terminating the 85-day maintenance cycle on the first two EMU (Extravehicular Mobility Unit) batteries in BCM3 (Battery Charger Module 3) and BCM4 [the periodic battery maintenance consists of fully discharging and then recharging the storage units to prolong their useful life];
  • Conducting the standard one-hour scrubs of EMUs 3005 & 3011 cooling loops with their SCUs (Service & Cooling Umbilicals), filtering ionic & particulate matter (via a 3-micron filter), then reconfiguring the cooling loops and starting the ~2hr biocide filtering [after scrubbing termination, Sandra disassembled the EMU water processing kit and performed SOP (Secondary Oxygen Pack) checkout & leak check on EMU 3005 (only). Loop scrubbing, incl. iodination of the LCVGs (Liquid Cooling & Ventilation Garments) for biocidal maintenance, is done to eliminate any biomass and particulate matter that may have accumulated in the loops. Maintenance on EMU 3004 was performed yesterday (3004 will return on 15A and would be used in the event of a Shuttle contingency EVA after undocking)]; and
  • Installing the LTA (Lower Torso Assembly) restraint bag for ease of transportation on EMU 3011(only) at the aft EDDA (EMU Don Doff Assembly).

The FE-2 also worked on the TOCA (Total Organic Carbon Analyzer), performing another calibration to check whether the previous calibration efforts, which she conducted with a Low & High TOC solution yesterday and on ¾, were successful. Results were transferred to SSC-7 (Station Support Computer 7) via USB drive for downlink and the data were also logged for calldown. [The re-calibration was necessary after the recent software change which decreased the “React TOC” oxidation reaction time setting from 10 min to 2 min to allow a one-run operation without exceeding the P4 sensor pressure limit (whereby H2/hydrogen and O2/oxygen are vented and dispersed into the cabin instead of utilizing TOCA’s catalytic action for elimination).]

On the FCF (Fluids & Combustion Facility) in the CIR (Combustion Integrated Rack) in the Lab, Sandy re-installed the alignment guides to lock down the PaRIS (Passive Rack Isolation System) after yesterday’s MDCA (Multi-user Droplet Combustion Apparatus)test session. [The ground conducted the first ignition of the CIR MDCA test yesterday, achieving three successful burns.]

In the US Lab, Fincke performed the regular controlled shut-down of the EHS VOA (Environmental Health System-Volatile Organic Analyzer), with the ground power-cycling its RPC-3 (Remote Power Controller 3), part of RPCM (RPC Module) LAD42B_A.

Magnus unstowed and set up the NUTRITION w/Repository hardware for a new session for urine and blood collections, starting tomorrow. For her blood draw tomorrow, Sandy’s 8-hr fast begins tonight.

Working on the GFI-11/OBSTANOVKA (Environment) equipment in the DC1, FE-1 Lonchakov installed a grounding strap on the Langmuir Probe to remove possible voltage during its use for electric field measurements on the external hull of the SM. [To date, field voltage measurements have been taken manually at selected times. Planning is underway to extend the electric measurements to continuous monitoring.]

Later, the FE-1 completed the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM. [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.]

Working from the discretionary “time permitting” task list, Yuri was also to perform the regular daily IMS (Inventory Management System) maintenance task by updating/editing the IMS standard “delta file” including stowage locations for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).

Sandra filled out the regular weekly FFQ (Food Frequency Questionnaire), her twelfth, on the MEC (Medical Equipment Computer). [On the FFQs, NASA astronauts keep a personalized log of their nutritional intake over time on special MEC software. Recorded are the amounts consumed during the past week of such food items as beverages, cereals, grains, eggs, breads, snacks, sweets, fruit, beans, soup, vegetables, dairy, fish, meat, chicken, sauces & spreads, and vitamins. The FFQ is performed once a week to estimate nutrient intake from the previous week and to give recommendations to ground specialists that help maintain optimal crew health. Weekly estimation has been verified to be reliable enough that nutrients do not need to be tracked daily.]

The station residents completed their regular daily 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the TVIS treadmill (CDR, FE-1/2.5h, FE-2) and ARED advanced resistive exerciser (CDR, FE-2).

Mike Fincke had three CDE (Crew Discretionary Event) audio/video downlinks scheduled, one at ~8:15am EST, the second at ~10:00am, the third at ~6:00pm, with the JAXA Flight Control Team, the Moscow Support Team, and the CEO (Crew Earth Observation) Group. Good thinking, Mike!

At ~5:20am, the crew held the regular (nominally weekly) tagup with the Russian Flight Control Team (GOGU), including Shift Flight Director (SRP), at TsUP via S-band/audio, phone-patched from Houston and Moscow.

At ~2:25pm, the crewmembers downlinked a PAO TV message of greetings to the veterans, alumni, faculty and students of the Military Training Institute at SibGAU (Academician M.F. Reshetnev Siberian State Aerospace University) at Krasnoyarsk on the occasion of its 45th Anniversary this month.

At ~4:30pm, the ISS crew will hold their regular weekly tagup with the Lead Flight Director at JSC/MCC-H via S-band/audio. [S/G-2 (Space-to-Ground 2) phone patch via SSC (Station Support Computer).]

At ~6:25pm, Sandy Magnus has a PFC (Private Family Conference) scheduled, via S-band/audio and Ku-band/MS-NetMeeting application (which displays the uplinked ground video on an SSC laptop).

WRM Update: A new WRM (Water Recovery Management) “cue card” was uplinked last night for the crew’s reference, updated with the latest water audit. [The new card (18-0006P) lists 37 CWCs (~1,081.0 L total) for the four types of water identified on board: technical water (640.3 L, for Elektron electrolysis, incl. 110.6 L currently off-limits, filled from WPA and pending sample analysis on the ground), potable water (390.1 L, incl. 174.6 L currently off-limit because of Wautersia bacteria), condensate water (0.0 L), waste/EMU dump and other (50.6 L, including the new CWC-I with 3.9 L from PWD flushes). Wautersia bacteria are typical water-borne microorganisms that have been seen previously in ISS water sources. These isolates pose no threat to human health.]

RS TVM Update: TsUP-Moscow this morning reported that yesterday’s attempt to restore the Russian TVM Terminal Computer System to full service on all three strings after the recent maintenance was not successful. Two of the three TVM lanes continue to function nominally, which is acceptable for the EVA-21A.

EVA-21A Timeline Preview: The Orlan EVA-21A by Lonchakov (EV1) & Fincke (EV2) on 3/10 begins at ~12:20pm EDT (DC1 EV hatch open), to last an estimated 5 hrs 45 min, i.e., concluding at approximately 5:05pm. Russian attitude thrusters will be inhibited by TsUP ground commanding at specific times when the spacewalkers work on the SM RO (Working Compartment, l.d.) and SM AO (Assembly Compartment). Objectives of the EVA (all fallen off the previous EVA-21 timeline) are –

  • Mount the EXPOSE-R hardware on the URM-D (Portable Multipurpose Workstation) on the SM RO l.d., connect it to the PF-3 connector patch panel and remove protective cover;
  • Photograph the URM-D with EXPOSE-R monoblock & cables, ROBOTIC hardware, IPI-SM hardware and routed cables;
  • Remove fasteners (Aramide straps) in the installation areas of the docking target and AR-VKA & 2AR-VKA antennas on DC1;
  • Close MLI (Multi-Layer Insulation) flap on the SM PF-10 connector patch panel;
  • Re-install SKK #9 removable cassette container in nominal position on SM;
  • Inspect & photograph Progress antenna ASF1-M-VKA from DC handrail 3034; and
  • Inspect & photograph conditions of ISS RS exterior & structural elements (“Panorama-2009” DTO).

CEO photo targets uplinked for today were Heard Island (Heard Island is a bleak, uninhabited, and mountainous island located in the Southern Ocean; about two-thirds of the way from Madagascar to Antarctica. Its mountains are covered in glaciers and dominated by Mawson Peak, a 9,006 ft high complex volcano which forms part of the Big Ben massif. A long thin spit named "Elephant Spit" extends from the east of the island. ISS pass was in mid-afternoon with at least partial clearing expected. Looking well right of track and trying for detailed oblique views of this challenging target), Patagonian Glaciers (ISS had a mid-afternoon, nadir pass with only partial clearing over the southern portion of the Southern Patagonian Ice Field. Specifically the CEO group is seeking nadir views of the Perito Moreno Glacier, a large [250 square kilometer in area with a length of 30km] feature on the southeastern flank of the ice field. It is one of three Patagonian glaciers that are not retreating. Mike & Sandy were asked to update CEO imagery of this beautiful glacier), andUlawun Volcano (Mount Ulawun is the highest volcano [2,334-m] of the Bismarck Archipelago and one of Papua New Guinea’s most frequently active. The crew had a mid-morning pass with partly cloudy conditions expected. ISS approach to the island of New Britain was from the NW with the giant island of New Guinea to the right. CEO has no useful photos of this volcano and is seeking only context views this time. There are a number of volcanoes on New Britain to the left of track. Ulawun is near the north coast across a strait from Lolobau Island).

CEO photography can be studied at this “Gateway” website: (as of 9/1/08, this database contained 770,668 views of the Earth from space, with 324,812 from the ISS alone).

ISS Orbit (as of this morning, 8:19am EST [= epoch])
Mean altitude — 355.1 km
Apogee height — 361.6 km
Perigee height — 348.7 km
Period — 91.64 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0009603
Solar Beta Angle — 33.9 deg (magnitude increasing)
Orbits per 24-hr. day — 15.71
Mean altitude loss in the last 24 hours — 57 m
Revolutions since FGB/Zarya launch (Nov. 98) — 58980

Significant Events Ahead (all dates Eastern Time, some changes possible!):
03/08/09 — US Daylight Time begins at 2:00am
03/08/09 — Russian EVA-21A suited dry run
03/10/09 — Russian EVA-21A (hatch open ~12:20pm EDT, 11:20am CDT)
03/11/09 — STS-119/Discovery/15A launch – S6 truss segment — 9:20:10pm EDT
03/13/09 — STS-119/Discovery/15A docking — 6:27pm EDT
03/23/09 — STS-119/Discovery/15A undocking — 10:23am EDT
03/25/09 — STS-119/Discovery/15A deorbit (Orbit 217) & landing — 3:27pm EDT (KSC)
03/26/09 — Soyuz TMA-14/18S launch
03/28/09 — STS-119/Discovery/15A landing (tentative)
03/28/09 — Soyuz TMA-14/18S docking (DC1)
04/07/09 — Soyuz TMA-13/17S undocking & landing
05/06/09 — Progress 32P undocking & deorbit (under review)
05/07/09 — Progress 33P launch
05/12/09 — STS-125/Atlantis Hubble Space Telescope Service Mission 4 (SM4)
05/12/09 — Progress 33P docking
05/15/09 — STS-127/Endeavour/2J/A launch – JEM EF, ELM-ES, ICC-VLD
05/27/09 — Soyuz TMA-15/19S launch
05/29/09 — Soyuz TMA-15/19S docking (FGB nadir)
Six-person crew on ISS
07/17/09 — Progress 33P undock & deorbit
07/20/09 — Soyuz TMA-14/18S relocation (to DC1)
07/24/09 — Progress 34P launch
07/26/09 — Progress 34P docking (SM aft)
08/06/09 — STS-128/Discovery/17A – MPLM (P), LMC
09/01/09 — H-IIB (JAXA HTV-1) — tentative
11/10/09 — Soyuz 5R/MRM2 (Russian Mini Research Module, MIM2) on Soyuz — tentative
11/12/09 — STS-129/Atlantis/ULF3 – ELC1, ELC2
12/10/09 — STS-130/Endeavour/20A – Node-3 + Cupola — tentative
02/11/10 — STS-131/Atlantis/19A – MPLM(P), LMC — tentative
04/08/10 — STS-132/Discovery/ULF4 – ICC-VLD, MRM1 — tentative
05/31/10 — STS-133/Endeavour/ULF5 – ELC3, ELC4 — tentative
12/XX/11 — Proton 3R/MLM w/ERA.

SpaceRef staff editor.